OpenStudy (anonymous):

Find the real or imaginary solutions of the equation by factoring: x^4 - 10x^2 + 16 = 0

Directrix (directrix):

x^4 - 10x^2 + 16 = 0 This will factor.

Directrix (directrix):

x^4 - 10x^2 + 16 = 0 (x^2 - 8) * (x^2 - 2) = 0 What comes next?